OpenStudy (anonymous):

This polynomial a perfect square. Which of these is its equivalent? x2 – 10x + 25

OpenStudy (anonymous):

well i don't see what answers you could choose from, but because you were given that it's a perfect square, immediately you can notice that in general, the problem will have this form\[(ax-b)^2\]where a and b are both constants, and for this problem, it comes out that if you factor it out this way, you'll get\[(x-5)^2=(x-5)(x-5)=x^2-10x+25\]